# Fuzzy dark matter simulations

### v1

@@ 2.3 Hybrid methods
Fluid-based schemes are significantly more efficient than wave-based schemes for addressing smooth, high-velocity flows, which occupy most of the simulation volume outside collapsed regions in cosmological simulations. Nevertheless, fluid schemes cannot robustly resolve strong interference in multi-stream regions, where wave schemes are preferred. This dilemma motivates hybrid approaches, which apply fluid schemes in smooth, single-stream regions and switch to wave schemes in interference-dominated, multi-stream regions. In addition, hybrid schemes facilitate cosmological zoom-in simulations, as purely wave-based schemes cannot accurately capture the large-scale structure when the de Broglie wavelength outside the zoom-in regions is poorly resolved. The Lagrangian zoom-in regions can be identified either by performing collisionless $N$-body simulations or by introducing tracer particles into FDM simulations, analogous to Bohmian mechanics (Wyatt 2005).
@@
Veltmaat et al. (2018) implemented a hybrid $N$-body–wave method in the AMR code Enzo. On coarser AMR levels (treated as fluid regions), the quantum pressure is ignored and the dynamics are approximated by a collisionless $N$-body approach. On the finest level, the code evolves the wave function using a finite-difference method. To reconstruct the wave function from $N$-body particles, a ‘classical wave function’ formulation is adopted, in which each particle represents a localized wave packet with a central phase $S$. The phase of particle $a$ is updated by solving Eq. (13) in Lagrangian form without the quantum potential $Q$:
@@
In these $N$-body–wave methods, the particles’ classical trajectories neglect the effects of quantum potential, and the reconstructed wave function is only accurate on a statistical level. To improve upon these limitations, Kunkel et al. (2025) proposed a fully grid-based hybrid scheme in the AMR code GAMER. On coarser levels, a finite-difference fluid scheme solving the Hamilton–Jacobi–Madelung equations in Eulerian form is employed (see Section 2.2.2). It retains the quantum potential term and thus remains accurate until divergent velocity and phase discontinuities develop at density voids. The phase difference between adjacent cells can substantially exceed $2\pi$ in general, reflecting the fact that fluid-based schemes do not need to resolve the de Broglie wavelength. Finer levels evolve the wave function using the local pseudo-spectral FC–Gram method described in Section 2.1.3. See Fig. 4 for an illustration.
Reconstructing the wave function from the fluid fields is unambiguous and straightforward, as both the density/phase and wave function are defined on grids, thus eliminating the need for an interpolation kernel. However, unlike in $N$-body–wave methods, the fluid scheme here also requires boundary conditions at the fluid–wave interfaces, since Eq. (13) involves spatial derivatives of both $S$ and $\rho$. A complication arises from the fact that the phase extracted from the wave function is only defined up to a multiple of $2\pi$. To resolve this ambiguity, the fine-grid phase in the wave region is unwrapped to align smoothly with the coarse-grid phase in the fluid region. This phase unwrapping implicitly assumes that the de Broglie wavelength is resolved by at least two grid points across the fluid–wave interfaces, which necessitates sufficiently fine resolution before switching to the wave formulation. While this may appear to increase computational cost, it is in fact a prerequisite for any wave-based schemes to avoid significant aliasing errors. See Section 2.4 for the corresponding grid refinement criteria.
@@ 2.5 Eigenmode methods
Here, $d\bm{p}$ is the discrete momentum-space resolution, and $N_{\bm{p}}$ are complex random numbers satisfying $\left\langle{N_{\bm{p}}^{*}N_{\bm{q}}}\right\rangle=\delta_{\bm{pq}}$, where $\left\langle{}\right\rangle$ denotes the ensemble average over different realizations. The distribution function $f(\bm{r},\bm{p})$ must satisfy the self-consistency condition: $m\sum_{\bm{p}}f(\bm{r},\bm{p})d\bm{p}=\rho_{0}(\bm{r})$. Under the assumptions of spherical symmetry and isotropy, $f(\bm{r},\bm{p})$ can be derived from the Eddington formula (Binney and Tremaine 2008; Teodori et al. 2025). To evolve the wave function using Eq. (56), the coefficients $a_{j}$ can be obtained by projecting the wave function from Eq. (59) onto the eigenmodes $\psi_{j}$. Unlike the method of Lin et al. (2018), this approach supports more general density profiles. One caveat, however, is that $f(\bm{r},\bm{p})$ cannot accurately model the central soliton core. This limitation may be addressed by superposing a theoretical soliton wave function following the soliton–halo relation. Finally, we note that Eq. (59) can also be used to construct a uniform FDM background with density granulation by inserting a Maxwell-Boltzmann distribution into $f(\bm{r},\bm{p})$ (Lancaster et al. 2020).
@@
This class of methods has been applied to a variety of studies, including the temporal and spatial correlation of density granules (Lin et al. 2018), soliton random walk, oscillations, and distortions (Li et al. 2021; Zagorac et al. 2022), and granulation-induced dynamical heating of stellar streams (Dalal et al. 2021) and ultra-faint dwarf galaxies (Dalal and Kravtsov 2022). In addition, eigenmode methods can be used solely to generate stable FDM initial conditions for Schrödinger–Poisson solvers. Example applications include studies of dynamical friction (Lancaster et al. 2020), soliton–halo relation (Yavetz et al. 2022), and granular heating of galactic disks (Yang et al. 2024) and dwarf spheroidal galaxies (Teodori et al. 2025).
2.6 Collisionless $N$-body methods
@@
The conventional collisionless $N$-body methods offer a convenient and efficient way to capture the large-scale dynamics of FDM, as the Schrödinger–Poisson equations reduce to the Vlasov–Poisson equations in the limit $\lambda_{\rm dB}\to 0$ (Widrow and Kaiser 1993). In practice, the simulation setup follows that of CDM simulations, except that the initial density power spectrum is generated using the FDM transfer function (Eqs. 19 and 20), making it analogous to warm dark matter (WDM) simulations. Compared to fully wave-based FDM simulations, this $N$-body approach allows for much larger comoving volumes due to its numerical efficiency. However, by neglecting quantum pressure, it fails to capture fine-grained wave phenomena, such as density granulation and soliton cores.
@@
The $N$-body methods, though approximate, have a number of practical applications. They can probe the FDM halo mass function in large-scale structure (Schive et al. 2016; May and Springel 2022), as the suppression of low-mass FDM halos is mainly determined by the simulation initial conditions (see Fig. 1). By comparing approximate $N$-body and genuine FDM simulations with the same initial conditions, one can quantify the dynamical effects of quantum pressure (Veltmaat and Niemeyer 2016; Nori and Baldi 2018; May and Springel 2021) and assess the numerical convergence of genuine FDM simulations (Liao et al. 2025, see also Fig. 16). In hybrid fluid–wave methods, the $N$-body approaches can provide the large-scale gravitational field and the boundary conditions for wave regions (see Section 2.3), and can also be used to identify the Lagrangian volumes of selected halos in cosmological zoom-in simulations (Veltmaat et al. 2018; Chan et al. 2025).
@@
Several studies have attempted to constrain the FDM particle mass using $N$-body simulations, such as through the Lyman-alpha forest power spectrum (e.g., Iršič et al. 2017) and the subhalo mass function (e.g., Nadler et al. 2025). However, it remains unclear whether omitting strong interference fringes and quantum pressure effects—such as their impact on the tidal evolution of subhalos and solitons (Chan et al. 2025)—may bias these constraints. Moreover, $N$-body approaches are susceptible to spurious halo formation, albeit with properties differing quantitatively from those in genuine FDM simulations (see Fig. 8).
@@ 3.2 Vortices
Solving the momentum conservation equation, Eq. (17), may seem promising at first glance, since at vortices the momentum density $\rho v$ vanishes, and both the momentum density flux $\rho v^{2}$ and the stress tensor $\Sigma_{ij}\propto\rho\frac{\partial^{2}\log\rho}{\partial x_{i}\partial x_{j}}$ remain finite (as shown in Fig. 14). However, computing the momentum density flux involves evaluating the ratio of two infinitesimal quantities, $(\rho v)^{2}$ and $\rho$. Likewise, an accurate estimation of the stress tensor depends on a delicate cancellation between the divergent term $\frac{\partial^{2}\log\rho}{\partial x_{i}\partial x_{j}}$ and the vanishing term $\rho$. These calculations are therefore highly susceptible to large numerical errors.
@@ 3.3 Spurious halos
Figure 8: An illustration of spurious halos in FDM simulations. The image compares the density distributions along a filament at $z=0$ across three different simulation setups from the same FDM initial condition: (top) a collisionless $N$-body simulation, (middle) a genuine FDM simulation with AMR using a hybrid fluid–wave algorithm, and (bottom) a genuine FDM simulation using the global Fourier method. Spurious halos, highlighted by arrows in the top and middle panels, appear along the filament as regularly spaced clumps with masses substantially below the half-mode mass, $M_{1/2}\propto m_{22}^{-4/3}$. Notably, these halos lack consistent counterparts in simulations with different setups, as they arise from artificial fragmentation seeded by numerical noise. The bottom panel shows no evidence of spurious halos, presumably due to the superior accuracy of the global Fourier method
Fig. 8 illustrates this issue by comparing the density distributions along a filament across three different simulation setups, all starting from the same FDM initial condition with $m_{22}=0.8$: (i) a collisionless $N$-body simulation (Section 2.6), (ii) a genuine FDM simulation using the hybrid fluid–wave algorithm based on the local pseudo-spectral FC–Gram method (Section 2.3), and (iii) a genuine FDM simulation using the global Fourier method (Section 2.1.1). The two massive halos at the ends of the filament are presumably genuine, as their masses are comparable to $M_{1/2}$ and they appear consistently in all three simulations. By contrast, the low-mass clumps along the filament in the upper and middle panels are likely artifacts, as they exhibit several features commonly associated with spurious halos. First, their masses are well below $M_{1/2}$. Second, their spatial distribution shows a suspiciously regular spacing along the filament. Third, these halos lack consistent counterparts in simulations with different setups, suggesting that they are sensitive to numerical accuracy.
Spurious halos in the $N$-body simulation are expected, since this setup is closely analogous to a WDM simulation, differing only in the exact shape of the initial density power spectrum. Their emergence in the genuine FDM simulation (middle panel) is also not surprising, given the above discussion, although this issue has received comparatively less attention than in the WDM case. For example, do spurious protohalos in FDM simulations also exhibit flatter geometries than genuine halos? How do their characteristic mass and spacing depend on resolution and algorithm? Notably, Fig. 8 suggests that the global Fourier method may be free from artificial fragmentation, possibly due to its superior accuracy along smooth filaments. However, this method cannot capture fine structures within the two massive halos once the de Broglie wavelength approaches the grid scale.
@@ 4.2.3 Isolated halos
As a demonstration, Fig. 13 shows the density and velocity profiles of an FDM halo extracted from a cosmological simulation with $m_{22}=0.8$ at $z=0$. We extract a cubic region of width $186\,\rm kpc$, centered on a halo with a virial radius $r_{\rm h}=72\,\rm kpc$ and a mass $M_{\rm h}=1.9\times 10^{10}\,\rm{M}_{\odot}$ (see Data availability for the link to download the halo data). The central soliton has a radius $r_{\rm s}=0.53\,\rm kpc$ and a mass $M_{\rm s}=1.5\times 10^{8}\,\rm{M}_{\odot}$, with a soliton–halo transition approximately at $4\,r_{\rm s}$. The halo density profile at $r\gtrsim 8\,r_{\rm s}$ is well fitted by the NFW model with a concentration parameter of $c=5.0$. To validate the system’s stability, we evolve it for another $2.3\,\rm Gyr$ (roughly one free-fall time), using the global Fourier method employed in the GAMER code on a $N=1024^{3}$ grid without AMR, corresponding to a spatial resolution of $0.18\,\rm kpc$. We apply isolated boundary conditions for the Poisson solver and periodic boundary conditions for the wave function.
@@
In addition to being extracted from cosmological simulations, isolated FDM halos can also be constructed from scratch via the gravitational collapse of a bound system (e.g., Li et al. 2021; Hui et al. 2021; Blum et al. 2025) or from a prescribed distribution function (see Section 2.5). Furthermore, beyond their use in calibrating FDM algorithms, isolated halos serve as versatile simulation testbeds for investigating several distinctive FDM phenomena, such as the formation and dynamics of quantized vortices (Hui et al. 2021), the stochastic motion of nuclear objects (Schive et al. 2020; Dutta Chowdhury et al. 2021; Li et al. 2021), and dynamical heating in dwarf galaxies (Dalal and Kravtsov 2022; Dutta Chowdhury et al. 2023; Teodori et al. 2025), galactic disks (Yang et al. 2024), and stellar streams (Dalal et al. 2021).

### v2

@@ 2.3 Hybrid methods
Fluid-based schemes are significantly more efficient than wave-based schemes for addressing smooth, high-velocity flows, which occupy most of the simulation volume outside collapsed regions in cosmological simulations. Nevertheless, fluid schemes cannot robustly resolve strong interference in multi-stream regions, where wave schemes are preferred. This dilemma motivates hybrid approaches, which apply fluid schemes in smooth, single-stream regions and switch to wave schemes in interference-dominated, multi-stream regions. In addition, hybrid schemes facilitate cosmological zoom-in simulations, as purely wave-based schemes cannot accurately capture the large-scale structure when the de Broglie wavelength outside the zoom-in regions is poorly resolved. The Lagrangian zoom-in regions can be identified either by performing collisionless NN-body simulations or by introducing tracer particles into FDM simulations, analogous to Bohmian mechanics (Wyatt 2005).
@@
Veltmaat et al. (2018) implemented a hybrid NN-body–wave method in the AMR code Enzo. On coarser AMR levels (treated as fluid regions), the quantum pressure is ignored and the dynamics are approximated by a collisionless NN-body approach. On the finest level, the code evolves the wave function using a finite-difference method. To reconstruct the wave function from NN-body particles, a ‘classical wave function’ formulation is adopted, in which each particle represents a localized wave packet with a central phase $S$. The phase of particle $a$ is updated by solving Eq. (13) in Lagrangian form without the quantum potential $Q$:
@@
In these NN-body–wave methods, the particles’ classical trajectories neglect the effects of quantum potential, and the reconstructed wave function is only accurate on a statistical level. To improve upon these limitations, Kunkel et al. (2025) proposed a fully grid-based hybrid scheme in the AMR code GAMER. On coarser levels, a finite-difference fluid scheme solving the Hamilton–Jacobi–Madelung equations in Eulerian form is employed (see Section 2.2.2). It retains the quantum potential term and thus remains accurate until divergent velocity and phase discontinuities develop at density voids. The phase difference between adjacent cells can substantially exceed $2\pi$ in general, reflecting the fact that fluid-based schemes do not need to resolve the de Broglie wavelength. Finer levels evolve the wave function using the local pseudo-spectral FC–Gram method described in Section 2.1.3. See Fig. 4 for an illustration.
Reconstructing the wave function from the fluid fields is unambiguous and straightforward, as both the density/phase and wave function are defined on grids, thus eliminating the need for an interpolation kernel. However, unlike in NN-body–wave methods, the fluid scheme here also requires boundary conditions at the fluid–wave interfaces, since Eq. (13) involves spatial derivatives of both $S$ and $\rho$. A complication arises from the fact that the phase extracted from the wave function is only defined up to a multiple of $2\pi$. To resolve this ambiguity, the fine-grid phase in the wave region is unwrapped to align smoothly with the coarse-grid phase in the fluid region. This phase unwrapping implicitly assumes that the de Broglie wavelength is resolved by at least two grid points across the fluid–wave interfaces, which necessitates sufficiently fine resolution before switching to the wave formulation. While this may appear to increase computational cost, it is in fact a prerequisite for any wave-based schemes to avoid significant aliasing errors. See Section 2.4 for the corresponding grid refinement criteria.
@@ 2.5 Eigenmode methods
Here, $d\bm{p}$ is the discrete momentum-space resolution, and $N_{\bm{p}}$ are complex random numbers satisfying $\left\langle{N_{\bm{p}}^{*}N_{\bm{q}}}\right\rangle=\delta_{\bm{pq}}$, where $\left\langle{}\right\rangle$ denotes the ensemble average over different realizations. The distribution function $f(\bm{r},\bm{p})$ must satisfy the self-consistency condition: $m\sum_{\bm{p}}f(\bm{r},\bm{p})d\bm{p}=\rho_{0}(\bm{r})$. Under the assumptions of spherical symmetry and isotropy, $f(\bm{r},\bm{p})$ can be derived from the Eddington formula (Binney and Tremaine 2008; Teodori et al. 2026). To evolve the wave function using Eq. (56), the coefficients $a_{j}$ can be obtained by projecting the wave function from Eq. (59) onto the eigenmodes $\psi_{j}$. Unlike the method of Lin et al. (2018), this approach supports more general density profiles. One caveat, however, is that $f(\bm{r},\bm{p})$ cannot accurately model the central soliton core. This limitation may be addressed by superposing a theoretical soliton wave function following the soliton–halo relation. Finally, we note that Eq. (59) can also be used to construct a uniform FDM background with density granulation by inserting a Maxwell-Boltzmann distribution into $f(\bm{r},\bm{p})$ (Lancaster et al. 2020).
@@
This class of methods has been applied to a variety of studies, including the temporal and spatial correlation of density granules (Lin et al. 2018), soliton random walk, oscillations, and distortions (Li et al. 2021; Zagorac et al. 2022), and granulation-induced dynamical heating of stellar streams (Dalal et al. 2021) and ultra-faint dwarf galaxies (Dalal and Kravtsov 2022). In addition, eigenmode methods can be used solely to generate stable FDM initial conditions for Schrödinger–Poisson solvers. Example applications include studies of dynamical friction (Lancaster et al. 2020), soliton–halo relation (Yavetz et al. 2022), and granular heating of galactic disks (Yang et al. 2024) and dwarf spheroidal galaxies (Teodori et al. 2026).
2.6 Collisionless NN-body methods
@@
The conventional collisionless NN-body methods offer a convenient and efficient way to capture the large-scale dynamics of FDM, as the Schrödinger–Poisson equations reduce to the Vlasov–Poisson equations in the limit $\lambda_{\rm dB}\to 0$ (Widrow and Kaiser 1993). In practice, the simulation setup follows that of CDM simulations, except that the initial density power spectrum is generated using the FDM transfer function (Eqs. 19 and 20), making it analogous to warm dark matter (WDM) simulations. Compared to fully wave-based FDM simulations, this NN-body approach allows for much larger comoving volumes due to its numerical efficiency. However, by neglecting quantum pressure, it fails to capture fine-grained wave phenomena, such as density granulation and soliton cores.
@@
The NN-body methods, though approximate, have a number of practical applications. They can probe the FDM halo mass function in large-scale structure (Schive et al. 2016; May and Springel 2022), as the suppression of low-mass FDM halos is mainly determined by the simulation initial conditions (see Fig. 1). They are also useful for studying the morphology of FDM halos (Dome et al. 2023). By comparing approximate NN-body and genuine FDM simulations with the same initial conditions, one can quantify the dynamical effects of quantum pressure (Veltmaat and Niemeyer 2016; Nori and Baldi 2018; May and Springel 2021) and assess the numerical convergence of genuine FDM simulations (Liao et al. 2025, see also Fig. 16). In hybrid fluid–wave methods, the NN-body approaches can provide the large-scale gravitational field and the boundary conditions for wave regions (see Section 2.3), and can also be used to identify the Lagrangian volumes of selected halos in cosmological zoom-in simulations (Veltmaat et al. 2018; Chan et al. 2025).
@@
Several studies have attempted to constrain the FDM particle mass using NN-body simulations, such as through the Lyman-alpha forest power spectrum (e.g., Iršič et al. 2017) and the subhalo mass function (e.g., Nadler et al. 2025). However, it remains unclear whether omitting strong interference fringes and quantum pressure effects—such as their impact on the tidal evolution of subhalos and solitons (Chan et al. 2025)—may bias these constraints. Moreover, NN-body approaches are susceptible to spurious halo formation, albeit with properties differing quantitatively from those in genuine FDM simulations (see Fig. 8).
@@ 3.2 Vortices
Solving the momentum conservation equation, Eq. (17), may seem promising at first glance, since at vortices the momentum density $\rho v$ vanishes, and both the momentum density flux $\rho v^{2}$ and the stress tensor $\Sigma_{ij}\propto\rho\frac{\partial^{2}\ln\rho}{\partial x_{i}\partial x_{j}}$ remain finite (as shown in Fig. 14). However, computing the momentum density flux involves evaluating the ratio of two infinitesimal quantities, $(\rho v)^{2}$ and $\rho$. Likewise, an accurate estimation of the stress tensor depends on a delicate cancellation between the divergent term $\frac{\partial^{2}\ln\rho}{\partial x_{i}\partial x_{j}}$ and the vanishing term $\rho$. These calculations are therefore highly susceptible to large numerical errors.
@@ 3.3 Spurious halos
Figure 8: An illustration of spurious halos in FDM simulations. The image compares the density distributions along a filament at $z=0$ across three different simulation setups from the same FDM initial condition: (top) a collisionless NN-body simulation, (middle) a genuine FDM simulation with AMR using a hybrid fluid–wave algorithm, and (bottom) a genuine FDM simulation using the global Fourier method. Spurious halos, highlighted by arrows in the top and middle panels, appear along the filament as regularly spaced clumps with masses substantially below the half-mode mass, $M_{1/2}\propto m_{22}^{-4/3}$. Notably, these halos lack consistent counterparts in simulations with different setups, as they arise from artificial fragmentation seeded by numerical noise. The bottom panel shows no evidence of spurious halos, presumably due to the superior accuracy of the global Fourier method
Fig. 8 illustrates this issue by comparing the density distributions along a filament across three different simulation setups, all starting from the same FDM initial condition with $m_{22}=0.8$: (i) a collisionless NN-body simulation (Section 2.6), (ii) a genuine FDM simulation using the hybrid fluid–wave algorithm based on the local pseudo-spectral FC–Gram method (Section 2.3), and (iii) a genuine FDM simulation using the global Fourier method (Section 2.1.1). The two massive halos at the ends of the filament are presumably genuine, as their masses are comparable to $M_{1/2}$ and they appear consistently in all three simulations. By contrast, the low-mass clumps along the filament in the upper and middle panels are likely artifacts, as they exhibit several features commonly associated with spurious halos. First, their masses are well below $M_{1/2}$. Second, their spatial distribution shows a suspiciously regular spacing along the filament. Third, these halos lack consistent counterparts in simulations with different setups, suggesting that they are sensitive to numerical accuracy.
Spurious halos in the NN-body simulation are expected, since this setup is closely analogous to a WDM simulation, differing only in the exact shape of the initial density power spectrum. Their emergence in the genuine FDM simulation (middle panel) is also not surprising, given the above discussion, although this issue has received comparatively less attention than in the WDM case. For example, do spurious protohalos in FDM simulations also exhibit flatter geometries than genuine halos? How do their characteristic mass and spacing depend on resolution and algorithm? Notably, Fig. 8 suggests that the global Fourier method may be free from artificial fragmentation, possibly due to its superior accuracy along smooth filaments. However, this method cannot capture fine structures within the two massive halos once the de Broglie wavelength approaches the grid scale.
@@ 4.2.3 Isolated halos
As a demonstration, Fig. 13 shows the density and velocity profiles of an FDM halo extracted from a cosmological simulation with $m_{22}=0.8$ at $z=0$. We extract a cubic region of width $186\,\rm kpc$, centered on a halo with a virial radius $r_{\rm h}=72\,\rm kpc$ and a mass $M_{\rm h}=1.9\times 10^{10}\,\rm{M}_{\odot}$ (see Data availability for the link to download the halo data). The central soliton has a radius $r_{\rm s}=0.52\,\rm kpc$ and a mass $M_{\rm s}=1.5\times 10^{8}\,\rm{M}_{\odot}$, with a soliton–halo transition approximately at $4\,r_{\rm s}$. The halo density profile at $r\gtrsim 8\,r_{\rm s}$ is well fitted by the NFW model with a concentration parameter of $c=5.4$. To validate the system’s stability, we evolve it for another $2.3\,\rm Gyr$ (roughly one free-fall time), using the global Fourier method employed in the GAMER code on a $N=1024^{3}$ grid without AMR, corresponding to a spatial resolution of $0.18\,\rm kpc$. We apply isolated boundary conditions for the Poisson solver and periodic boundary conditions for the wave function.
@@
In addition to being extracted from cosmological simulations, isolated FDM halos can also be constructed from scratch via the gravitational collapse of a bound system (e.g., Li et al. 2021; Hui et al. 2021; Blum et al. 2025) or from a prescribed distribution function (see Section 2.5). Furthermore, beyond their use in calibrating FDM algorithms, isolated halos serve as versatile simulation testbeds for investigating several distinctive FDM phenomena, such as the formation and dynamics of quantized vortices (Hui et al. 2021), the stochastic motion of nuclear objects (Schive et al. 2020; Dutta Chowdhury et al. 2021; Li et al. 2021), and dynamical heating in dwarf galaxies (Dalal and Kravtsov 2022; Dutta Chowdhury et al. 2023; Teodori et al. 2026), galactic disks (Yang et al. 2024), and stellar streams (Dalal et al. 2021).